OpenStudy (anonymous):

Write two sets of parametric equations for .....y=x^4-1 @jim_thompson5910

OpenStudy (anonymous):

I used 20t

jimthompson5910 (jim_thompson5910):

One set I can think of off the top of my head is this x = t y = t^4-1 notice how because x = t, we can effectively replace every copy of 't' with x in the second equation

OpenStudy (anonymous):

i did 20t=x^4-1

jimthompson5910 (jim_thompson5910):

I'm not sure how you got 20t=x^4-1

OpenStudy (anonymous):

instead of using t...im using 20t

OpenStudy (anonymous):

i replaced y with 20t

OpenStudy (anonymous):

this isnt the answer, its the first thing i set up

jimthompson5910 (jim_thompson5910):

do you see how I got x = t y = t^4-1

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

that's one set. I'm trying to think of the other

jimthompson5910 (jim_thompson5910):

let's solve y = t^4-1 for t y = t^4-1 y+1 = t^4 t^4 = y+1 t = (y+1)^(1/4) if y = t, then x = (y+1)^(1/4) = (t+1)^(1/4) to ensure that x is both positive and negative, we use plus/minus, so the other set would be \[\Large \begin{cases}x = \pm(t+1)^{1/4} \\ {} \\ y = t\end{cases}\]

OpenStudy (anonymous):

ty

jimthompson5910 (jim_thompson5910):

no problem